7y-x=32
x=3y
Solve by using the substitution method.

Please help me with this if you can. Thank you.

Answer by stanbon(75887)

7y-x=32
x=3y
Solve by using the substitution method.
-------------------
Substitute for "x" and solve for "y":
7y -3y = 32
4y = 32
y = 8
----
substitute and solve for "x":
x = 3(8)
x = 24
================================
Cheers,
Stan H.